OpenStudy (anonymous):

Three Variable Systems of Equations Find the value of y in the solution to the system of equations. 2x + 2y + 3z = 3 x + 2y + 2z = -1 x - 3y + 5z = 6

OpenStudy (anonymous):

2x + 2y + 3z = 3 x + 2y + 2z = -1 -->(-2) -------------- 2x + 2y + 3z = 3 -2x - 4y - 4z = 2 (result of multiplying by -2) --------------add -2y - z = 5 x + 2y + 2z = -1 -->(-1) x - 3y + 5z = 6 --------------- -x - 2y - 2z = 1 (result of multiplying by -1) x - 3y + 5z = 6 --------------add -5y + 3z = 7 -2y - z = 5 -->(3) -5y + 3z = 7 ------------ -6y - 3z = 15 (result of multiplying by 3) -5y + 3z = 7 ------------add -11y = 22 y = -2 -5y + 3z = 7 -5(-2) + 3z = 7 10 + 3z = 7 3z = 7 - 10 3z = -3 z = -1 2x + 2y + 3z = 3 2x + 2(-2) + 3(-1) = 3 2x - 4 - 3 = 3 2x - 7 = 3 2x = 3 + 7 2x = 10 x = 5 check... x + 2y + 2z = -1 5 + 2(-2) + 2(-1) = -1 5 - 4 - 2 = -1 5 - 6 = -1 -1 = -1 (correct) x = 5, y = -2, z = -1

OpenStudy (anonymous):

Thanks a bunch!